OpenStudy (mindblast3r):

can a fraction be a a decimal number for example 1.5/2?

OpenStudy (anonymous):

No, you would have to simplify it.

OpenStudy (mindblast3r):

can a ratio be 1.5:2

OpenStudy (anonymous):

Yes

OpenStudy (mindblast3r):

how to simplify it now?

OpenStudy (anonymous):

1.5 = 3/2

OpenStudy (anonymous):

\[\frac{\frac{3}{2}}{2} = \frac{3}{2\times2}=\frac{3}{4}\]

OpenStudy (mindblast3r):

who wants medal?

OpenStudy (mindblast3r):

medal for all :D

OpenStudy (mindblast3r):

what the heck? lol i awarded you one o.O

OpenStudy (anonymous):

I'll take one from mind since he already had 3 lol, now everyones happy.

OpenStudy (anonymous):

>_>

OpenStudy (mindblast3r):

yay :D

OpenStudy (anonymous):

Now mind doesn't have one :c

OpenStudy (mindblast3r):

lol stop dis

OpenStudy (anonymous):

This is more complicated than any math problem.

OpenStudy (mindblast3r):

alright going to help other people with questions now :P